OpenStudy (anonymous):

x 5 ___ = ____ x + 9 8 find x

OpenStudy (anonymous):

Multiply the equation by the LCD: 8(X+9)

OpenStudy (diyadiya):

\[\frac{x}{x+9}= \frac{5}{8}\] cross-multiply Now we have \[8x= 5(x+9)\]\[8x=5x+45\] Now can you find x?

OpenStudy (anonymous):

That works too :P

OpenStudy (anonymous):

no i need x lol

OpenStudy (anonymous):

i dont know how to cross multiplyyyy

OpenStudy (diyadiya):